OpenStudy (anonymous):

hhhheeeellpppp! Calculus! Evaluate the derivative at of the function at the indicated point. f(x)=1/(x^2-3x)^2 (4, 1/16)

OpenStudy (raden):

use chain rule let u=x^2-3x du/dx= .....

OpenStudy (anonymous):

okay thanks!

OpenStudy (raden):

next, y=u^2 dy/du= ....... multiply of both to get dy/dx

OpenStudy (anonymous):

hello! you got this?

OpenStudy (anonymous):

not REALLY

OpenStudy (anonymous):

HELOO!

OpenStudy (anonymous):

ok as usual we go slow if you had\[\frac{1}{x^2}\] you would know what to do, right?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

you mean "yes!" or "yes?"

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

ok so i will take that as a maybe if you just had \(\frac{1}{x^2}\) you could use the power rule, after writing \[\frac{1}{x^2}=x^{-2}\]

OpenStudy (anonymous):

yes!!!!

OpenStudy (anonymous):

then by the power rule, the derivative would be ...

OpenStudy (anonymous):

-2x^-3

OpenStudy (anonymous):

right, also known as \(-\frac{2}{x^3}\)

OpenStudy (anonymous):

yes :}

OpenStudy (anonymous):

BUT you do not have \(\frac{1}{x^2}\) you have \[\frac{1}{(x^2-3x)^2}\] so you have to adjust

OpenStudy (anonymous):

kay:l

OpenStudy (anonymous):

so here is what you do take your previous answer, \(-\frac{2}{x^3}\) and replace \(x\) by \(x^2-3x\) so instead of getting \[-\frac{2}{x^3}\] you get \[\frac{-2}{(x^2-3x)^3}\]

OpenStudy (anonymous):

THEN multiply this by the derivative of \(x^2-3x\)

OpenStudy (anonymous):

let me know if you have a question

OpenStudy (anonymous):

u confused the s%&# outta me boy or girl satellite 73

OpenStudy (anonymous):

ho ho ho i see i am good at that

OpenStudy (anonymous):

thanks so much encouragement :p

OpenStudy (anonymous):

what is the derivative of \(\sin(x)\) ?

OpenStudy (anonymous):

cosx

OpenStudy (anonymous):

ok good now what is the derivative of \(\sin(x^2)\) ?

OpenStudy (anonymous):

man ummmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmmm cosx^2????

OpenStudy (anonymous):

hmmmmmmmmmmmmmmmmm nope

OpenStudy (anonymous):

you have to take \(\cos(x^2)\) and multiply it by the derivative of \(x^2\) which is \(2x\)

OpenStudy (anonymous):

this is called the "chain rule" because you have a chain of relations or functions. we say for \(\sin(x^2)\) the "inside" function is \(x^2\) and the "outside" function is \(\sin\)

OpenStudy (anonymous):

so you have a composition of functions

OpenStudy (anonymous):

so it'll be 2xcosx?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

YEAH!!!!! yaee!

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

don't change the inside piece, it is still \(x^2\)

OpenStudy (anonymous):

o okay 2xcosx^2?

OpenStudy (anonymous):

the answer would be \(2x\cos(x^2)\)

OpenStudy (anonymous):

right in english: the derivative of a composite function is: the derivative of the "outside" function, evaluated at the inside function, times the derivative of the "inside" function

OpenStudy (anonymous):

so for example, since the derivative of sine is cosine, then the derivative of \(\sin(x^2)\) is \[2x\cos(x^2)\]

OpenStudy (anonymous):

how about the derivative of \((3x^2-x)^{10}\) ?

OpenStudy (anonymous):

here the "inside" function is \(3x^2-x\) while the "outside" function is "to the tenth power""

OpenStudy (anonymous):

its aaaaaaaaaaaaaaa........................ hmmmmmmmmmummmmmmmmmmmmmmmmmmm.............. 10(3x^2-x) 10(6x-1)=======60x-1! :}

OpenStudy (anonymous):

NO! 60x-10

OpenStudy (anonymous):

:]

OpenStudy (anonymous):

close, very close it is \[10(3x^2-x)^9\times (6x-1)\]

OpenStudy (anonymous):

you forgot that the derivative of something to the power of ten is ten something to the power of nine by the power rule

OpenStudy (anonymous):

cuz its u*u(prime) right?

OpenStudy (anonymous):

mmk

OpenStudy (anonymous):

you forgot the 9th power

OpenStudy (anonymous):

lets try this \[(5x^2-3)^6\]

OpenStudy (anonymous):

last one cuz i gotta go satellite 73 (sorry :[ ) 6(5x^2-3)^5=========6(10x-3)^5======(60x-18)^5?????

OpenStudy (anonymous):

it is \[6(5x^2-3)^5\times (10x)\]

OpenStudy (anonymous):

you cannot distribute because of that exponent

OpenStudy (anonymous):

but that is the idea

OpenStudy (anonymous):

so you have \[(x^2-3x)^{-2}\] derivative is \[-2(x^2-3x)^{-3}\times (2x-3)\]

OpenStudy (anonymous):

ok :]

OpenStudy (anonymous):

bye~

OpenStudy (anonymous):

good luck bye